Same company as in RA 5.9: Stock price of $42, earnings of $2.12 per share during the last twelve months, forecasted earnings of $2.84 over the following year, and average earnings growth forecast of 12.5% per year for the next five years. What is this stock's PEG, rounded to one decimal place?
Answer #1

PEG using leading EPS=42/(2.84*12.5)=1.2
PEG using trailing EPS=42/(2.12*12.5)=1.6